OpenStudy (anonymous):

kendrick travels 2/3 mile to his friends house.He cycyles 3/4 of this distance and walks the rest of the way.What fraction of a mile did he cycle?

OpenStudy (anonymous):

1/6 of a mile (I'll show steps if you need them)

OpenStudy (anonymous):

oops...should have read the whole question (he walks 1/6 of a mile), but he BIKES 1/2 of a mile (3/4) * (2/3) = 6/12 = 1/2

OpenStudy (anonymous):

i don't get it